Hi! I can't figure this out, can you please help me?

The Question is:

Explain the following:

Nitrogen dioxide dissociates in the presence of solar radiation to produce free radicals of oxygen. Explain.

Thank you for helping me!
1
Expert's answer
2013-01-28T10:30:38-0500
Molecule of NO2 needs some energy for breaking chemical bond betweenO and N , and this energy can be given by sun.The most common radical in the lower atmosphere is molecular dioxygen. Photodissociation of source molecules produces other free radicals. In the lower atmosphere, the most important examples of free radical production are the photodissociation of nitrogen dioxide to give an oxygen atom and nitric oxide, which plays a key role in smog formation—and the photodissociation of ozone to give the excited oxygen atom O(1D). The chemical reactions for this process are:

R1 = NO2 + hν → NO + O
R2 = O + O2 → O3
R1 + R2 = NO2 + O2 + hν → NO + O3 <-- Formation ozone at ground level
R3 = NO + O3 → NO2 + O2 <-- Destruction of ozone